OpenStudy (anonymous):

2700=1350(1+(.09/1))^t Find the value of t using a logarithmic function, please show work and explain

OpenStudy (anonymous):

t=log((.09/1),1)

OpenStudy (anonymous):

just simplify the equation and then use a calculator

OpenStudy (anonymous):

How does the comma work?

OpenStudy (anonymous):

Can you do it and tell me exactly what you typed in please?

OpenStudy (dumbcow):

(1.09)^t = 2 take log of both sides log(1.09)^t = log (2) log property allows you to take exponent and bring it in front of log -> t*log(1.09) = log(2) t = log(2)/log(1.09)

OpenStudy (anonymous):

i just type \[\log_{a} b\] into my calc, it is the same as log(a,b)